# TAR3D: Creating High-Quality 3D Assets via Next-Part Prediction

### v2

@@ Abstract
footnotetext: *Equal contribution.          Corresponding author. footnotetext: §Work done during the internship at Shanghai AI Lab.
@@ 2 Related Work
Early 3D generation approaches mainly focus on the generation of different forms of 3D models, e.g., point clouds [67], meshes [9], and volumes [5] in a text/image-conditioned or unconditional manner. Limited by the categories and quantity of 3D objects used for training, these methods often do not generalize well. With the remarkable progress achieved by diffusion models in 2D image generation, a large number of researchers have explored migrating the pre-trained 2D priors to 3D generation. Pioneer works like DreamFusion [48], SJC [62], and Fantasia3D [7] feed the rendered views to a pre-trained 2D diffusion model and perform per-shape optimization for knowledge distillation. Despite ushering in a new era, these methods suffer from a series of serious issues, e.g., time-consuming and multi-face. Another line of methods like Zero123 [38], Zero123++ [54], and One2345 [36] use the rendered views of 3D objects to finetune the pre-trained diffusion models for new-view or multi-view generation. To improve the consistency between the generated multiple views, Consistent123 [35] and Cascade-Zero123 [10] introduce extra priors, e.g., boundary, and redundant views. SyncDreamer [39] proposes to correlate the corresponding features across different views by building a 3D-aware attention mechanism. These methods can be followed by a sparse-view reconstruction model, e.g., NeRFs [45, 63, 47], LRMs [27, 24, 70] and LGM [59], to generate 3D objects. However, the indirect generation fashion of these methods may lead to detail loss or reconstruction failures due to their heavy reliance on the fidelity of multi-view images.
More recently, a surge of works [74, 78, 68] directly synthesize 3D objects via a 3D-aware diffusion model, in which the 3D shapes are fed to a pre-trained 3D VAE for continuous latent features. Unlike these methods, our TAR3D eschews the diffusion scheme and creates 3D objects by autoregressively generating discrete geometric parts.
@@ 2.2 VAE & VQ-VAE
Variational Autoencoder (VAE) [31] is usually utilized to map high-dimensional input information to continuous probabilistic latent representations and has a far-reaching impact in the field of generative modeling. Thanks to this work, recent diffusion models [26, 52, 53] can be trained on limited computational resources while retaining their quality and flexibility. Recent 3D generation approaches like Clay [76], Direct3D [68], and LN3Diff [32] also explore to construct 3D-aware diffusion models based on this technology. Vector Quantised-Variational AutoEncoder (VQ-VAE) [61] is a variant of VAE. It introduces the codebook mechanism to quantize the continuous latent representations into discrete components, achieving promising performance on many generative tasks, e.g., text-to-image generation [51], music generation [19, 18], and speech gesture generation [2]. VQ-GAN [21] proposes to promote VQ-VAE by incorporating adversarial loss during the training process. In this paper, we build a 3D-aware VQ-VAE to quantize the triplane features of 3D shapes and acquire discrete geometric parts for autoregressive 3D generation.
@@ 3 Methodology
$z_{q}:=\bigg{(}\mathop{\mathrm{missing}}{arg~{}min}_{z_{k}\in\mathcal{Z}}||\tilde{\mathrm{z}}_{ij}-z_{k}||\bigg{)}\in\mathbb{R}^{B\times(3\times h\times w)\times d_{q}}$ (1)
@@ 3.3 Optimization Details
where $\hat{\mathcal{O}}(\cdot)$, and $\mathcal{O}(\cdot)$ are the predicted occupancy value and ground-truth occupancy value of the query point. For the codebook learning of quantizer, the training loss can be formulated to minimize the difference between the original features and the quantified features:
$\mathcal{L}_{cb}=||sg(\tilde{\mathrm{z}})-z_{q}||_{2}^{2}+\beta||\tilde{\mathrm{z}}-sg[z_{q}]||,$ (4)
@@ 4.1 Experiment Setup
Implementation details. In our 3D VQ-VAE, the number of point clouds input to the 3D shape encoder, i.e., $N_{p}$, is 81,920. The number of the self-attention layer in the 3D shape encoder, i.e., $N_{1}$, is 8. The height and width of the triplane feature maps, i.e., $h$ and $w$, are both set to 32. The channel numbers of the learnable tokens, and the triplane features, i.e., $d_{e}$ and $d_{z}$, are set to 768 and 16. The size and channel number of the codebook, i.e., $K$ and $d_{q}$, are set to 16,384 and 8 respectively. Besides, the number of the attention layers in the 3D geometry decoder is set as $N_{2}\mathrm{=}6$. The triplane features are upsampled to a resolution of $256\mathrm{\times}256$, in which 20,480 uniformly sampled voxel points and 20,480 near-surface points are employed as query points for occupancy supervision. The hyperparameters in Eqn. (5) are set as $\lambda_{rec}\mathrm{=}1$ and $\lambda_{cb}\mathrm{=}0.1$. We employ the CosineAnnealing scheduler [41], where the learning rate is initialized to $41e\mathrm{-}$ 4 and gradually decays over time. We adopt the AdamW optimizer [42] to train the 3D VQ-VAE with a total batch size of 128 for 100K steps on 8 NVIDIA A100 GPUs.
In our 3D GPT, we adopt the pre-trained DINO [75](ViT-B16) and FLAN-T5 XL [14] to encode the conditional images and text prompts respectively. As far as the decoder-only transformer, we follow the GPT-L setting of LLamaGen [57], which consists of 24 transformer layers with a head number of 16 and a dimension of 1024. We employ the AdamW optimizer with a learning rate of $41e\mathrm{-}$ 4 and a total batch size of 80 to train our 3D GPT for 100K steps. In addition, the classifier-free guidance (CFG) [25] scale of 7.5 is also introduced in the autoregressive inference to improve geometry quality and image/text-3D alignment.
@@ 4.3 Quantitative Evaluation
In this subsection, we use the mixed evaluation set of ShapeNet and Objaverse to quantitatively measure the performance of our TAR3D model against recent 3D object generation methods in terms of 2D visual quality and 3D geometric quality. Considering the diversity of 3D objects generated from text descriptions, we conduct experiments on image-to-3D tasks to ensure an accurate comparison, as shown in Tab. 1,. In particular, all candidate approaches employ the same images as their conditional input to generate 3D meshes. For the 2D evaluation, we render 20 views with $224\times 224$ resolution for each mesh and compare the yielding normal maps with the ones of corresponding ground-truth views. For the 3D evaluation, we compare the generated meshes with the ground truth meshes by uniformly sampling 16K point clouds from their surfaces in an aligned cube coordinate system of $[-1,1]^{3}$. Our TAR3D model surpasses other cutting-edge 3D object generation methods across all 2D and 3D metrics by a large margin. These experimental results further demonstrate the superiority of our TAR3D over existing methods.
@@
Triplane Size $3\!\times\!16\!\times\!16$ $3\!\times\!32\!\times\!32$ $3\!\times\!48\!\times\!48$
@@ 4.4 Ablation Studies
Triplane Size. We ablate the triplane size with 3$\times$16$\times$16, 3$\times$32$\times$32, and 3$\times$48$\times$48 to explore the impact of sequence length on the performance and efficiency. Note that the 3$\times$ 32$\times$32 widely adopted in 3D reconstruction is our default setting. As shown in Tab. 4, it is the best setting for efficiency and performance trade-offs, which is also demonstrated by the visual comparisons in Fig. 7
@@ 5 Conclusions & Future Work
In this paper, we propose a novel framework named TAR3D to generate high-quality 3D assets via the “next-token prediction” paradigm derived from multimodal large language models. To achieve this, we first develop a 3D VQ-VAE, in which the 3D shapes are encoded into the triplane latent space and quantized as discrete embeddings from a trainable codebook. Equipped with the well-designed feature deformation and attention mechanism, these quantized features are utilized to reconstruct fine-grained geometries in a neural occupancy field. Then, we adopt the codebook indices of these discrete representations to form the sequence for autoregressive modeling. Driven by the prefilling prompt embeddings, our 3D GPT, along with the 3D spatial information from the proposed TriPE, can generate high-quality 3D objects part by part. Finally, extensive experiments on several benchmark and out-of-domain 3D datasets, are conducted to demonstrate the superior performance of our TAR3D compared to existing approaches.
@@
We believe it is promising to improve 3D object generation under the autoregressive setting. There are a few foreseeable directions to explore: 1) Scaling law. We will collect more 3D data with prompts from other datasets, e.g., Objaverse-XL [17], 3D-FUTURE [22], and scale the GPT model for a stronger generation capability. 2) Sequence formulation. We will explore the sequence formulation manner of triplane indices for more efficient generative modeling. 3) Tokenization. We plan to establish a unified codebook for 3D representations and prompt embeddings.
[…]
Supplementary Material
Appendix A More Visualization Results
Figure 8: Given a wide range of input images from the benchmark datasets, i.e., ShapeNet and Objaverse, our TAR3D model is capable of inferring the full picture of the primary objects and producing high-quality 3D meshes that adhere to them.
Figure 9: Given a wide range of input images from the out-of-domain dataset, i.e., GSO, our TAR3D model is capable of inferring the full picture of the primary objects and producing high-quality 3D meshes that adhere to them.
Figure 10: Qualitative results of our TAR3D on a broad spectrum of text prompts. It can be observed that these 3D meshes created by our TAR3D model are semantically faithful to the textual descriptions.
Apart from the quantitative and qualitative comparisons of the proposed TAR3D with recent representative conditional 3D object generation approaches in the main text, in this section, we also provide more synthesized 3D meshes of TAR3D under image-to-3d and text-to-3d settings.
Fig. 8 and Fig. 9 show the generated meshes when feeding our TAR3D with image prompts from the benchmark datasets (Objaverse and ShapeNet) and the out-of-domain dataset (GSO), respectively. Our TAR3D can infer the full picture of an object from specific views, resulting in 3D assets that adhere to the given images.
Fig. 10 illustrates the 3D objects created by our TAR3D according to various textual descriptions. Note that these text prompts come from the ULIP dataset [71], in which the texts are utilized to describe the rendered images while containing some noise for 3D objects, e.g., “a grey background” in the sentence describing the camera model. Nevertheless, our TAR3D can still generate faithful 3D assets for the primary objects described in these sentences.
Appendix B More Discussions about Future Work
Although this work establishes an autoregressive 3D generation pipeline based on triplane representations from scratch, it still has certain limitations. We summarize them as follows and leave them for future research.
  • Data Cleaning and Expansion. The prompt system of our TAR3D is built on the ULIP dataset [71]. However, the quality of the rendered images is unsatisfactory, and there is a certain level of noise in the textual descriptions. We will re-render the collected 3D objects and adopt the Cap3D strategy [44] to caption them. In addition, we will collect more data with prompts from other 3D datasets, like Objaverse-XL [17], Omniobject3d [69], ABO [15], etc.
  • Model Scaling Up. The 3D GPT architecture is almost the same as Llama. Once the aforementioned 3D dataset is assembled, it is promising to scale the GPT model via optimization techniques and training recipes in the LLM community for a stronger 3D generation capability.
  • Sequence formulation. The sequence length in this work is three times that of image autoregressive generation, as the triplane representation consists of three feature maps. Although shorter than methods based on mesh face quantization, it still has room for improvement. In future work, we will explore more concise sequence formulations of triplane indices for efficient autoregressive modeling.
  • Tokenization. In this paper, we employ the encoded image or text features as prefilling token embeddings to integrate such conditions into autoregressive models. Considering the powerful multimodal unification capability of the GPT model, we will establish a unified codebook for 3D representations and prompt embeddings.

### v3

@@ Abstract
00footnotetext: *Equal contribution.          Corresponding author. 00footnotetext: §Work done during the internship at Shanghai AI Lab.
@@ 2 Related Work
Early 3D generation approaches mainly focus on the generation of different forms of 3D models, e.g., point clouds [71, 30, 82], meshes [9, 74, 87, 88], and volumes [5, 66, 58] in a text/image-conditioned or unconditional manner. Limited by the categories and quantity of 3D objects used for training, these methods often do not generalize well. With the remarkable progress achieved by diffusion models in 2D image generation, a large number of researchers have explored migrating the pre-trained 2D priors to 3D generation. Pioneer works [49, 65, 7, 84] feed the rendered views to a 2D pre-trained model and perform per-shape optimization for knowledge distillation. Despite ushering in a new era, these methods suffer from a series of serious issues, e.g., time-consuming and multi-face. Another line of methods like Zero123/++ [40, 55], One2345 [38], and AR123 [85] use the rendered views of 3D objects to finetune the pre-trained diffusion models for new-view or multi-view generation. To improve the consistency between the generated multiple views, Consistent123 [37] and Cascade-Zero123 [10] introduce extra priors, e.g., boundary, and redundant views. SyncDreamer [41] proposes to correlate the corresponding features across different views by building a 3D-aware attention mechanism. These methods can be followed by a sparse-view reconstruction model, e.g., NeRFs [46, 67, 48], LRMs [27, 25, 73] and LGM [62], to generate 3D objects. However, the indirect generation fashion of these methods may lead to detail loss or reconstruction failures due to their heavy reliance on the fidelity of multi-view images.
More recently, a surge of works [78, 86, 72, 14] directly synthesize 3D objects via a 3D diffusion model, in which the 3D shapes are fed to a pre-trained 3D VAE for continuous latent features. Unlike these methods, our TAR3D eschews the diffusion scheme and creates 3D objects by autoregressively generating discrete geometric parts.
@@ 2.2 VAE & VQ-VAE
Variational Autoencoder (VAE) [32] is usually utilized to map high-dimensional input information to continuous probabilistic latent representations and has a far-reaching impact in the field of generative modeling. Thanks to this work, recent diffusion models [53, 54, 12, 60] can be trained on limited computational resources while retaining their quality and flexibility. Recent 3D generation approaches like Clay [80], Direct3D [72], and LN3Diff [33] also explore to construct 3D-aware diffusion models based on this technology. Vector Quantised-Variational AutoEncoder (VQ-VAE) [64] is a variant of VAE. It introduces the codebook mechanism to quantize the continuous latent representations into discrete components, achieving promising performance on many generative tasks, e.g., text-to-image generation [52], music generation [20, 19], and speech gesture generation [2]. VQ-GAN [22] proposes to promote VQ-VAE by incorporating adversarial loss during the training process. In this paper, we build a 3D-aware VQ-VAE to quantize the triplane features of 3D shapes and acquire discrete geometric parts for autoregressive 3D generation.
@@ 3 Methodology
[…]
@@ 3.3 Optimization Details
[…]
@@ 4.1 Experiment Setup
Implementation details. In our 3D VQ-VAE, the number of point clouds input to the 3D shape encoder, i.e., $N_{p}$, is 81,920. The number of the self-attention layer in the 3D shape encoder, i.e., $N_{1}$ is 8. The height and width of the triplane feature maps, i.e., $h$ and $w$, are both set to 32. The channel numbers of the learnable tokens, and the triplane features, i.e., $d_{e}$ and $d_{z}$, are set to 768 and 16. The size and channel number of the codebook, i.e., $K$ and $d_{q}$, are set to 16,384 and 8, respectively. Moreover, the number of the attention layers in the 3D geometry decoder is set as $N_{2}\mathrm{=}6$. The triplane features are upsampled to a resolution of $256\mathrm{\times}256$, in which 20,480 uniformly sampled voxel points and 20,480 near-surface points are employed as query points for occupancy supervision. The hyperparameters in Eqn. (5) are set as $\lambda_{rec}\mathrm{=}1$ and $\lambda_{cb}\mathrm{=}0.1$. We employ the CosineAnnealing scheduler [43], where the learning rate is initialized to $1e\mathrm{-}4$ and gradually decays over time. We adopt the AdamW optimizer [44] to train the 3D VQ-VAE with a total batch size of 128 for 100K steps on 8 NVIDIA A100 GPUs.
In our 3D GPT, we adopt the pre-trained DINO [79](ViT-B16) and FLAN-T5 XL [16] to encode the conditional images and text prompts respectively. As far as the decoder-only transformer, we follow the GPT-L setting of LLamaGen [59], which consists of 24 transformer layers with a head number of 16 and a dimension of 1024. We employ the AdamW optimizer with a learning rate of $1e\mathrm{-}4$ and a total batch size of 80 to train our 3D GPT for 100K steps. In addition, the classifier-free guidance (CFG) [26] scale of 7.5 is also introduced in the autoregressive inference to improve geometry quality and image/text-3D alignment.
@@ 4.3 Quantitative Evaluation
In this subsection, we use the mixed evaluation set of ShapeNet and Objaverse to quantitatively measure the performance of our TAR3D model against recent 3D object generation methods in terms of 2D visual quality and 3D geometric quality. Considering the diversity of 3D objects generated from text descriptions, we conduct experiments on image-to-3D tasks to ensure an accurate comparison, as shown in Tab. 1,. In particular, all candidate approaches employ the same images as their conditional input to generate 3D meshes. For the 2D evaluation, we render 20 views with $224\times 224$ resolution for each mesh and compare the yielding normal maps with the ones of corresponding ground-truth views. For the 3D evaluation, we compare the generated meshes with the ground truth meshes by uniformly sampling 16K point clouds from their surfaces in an aligned cube coordinate system of $1,1]^{3}$. Our TAR3D model surpasses other cutting-edge 3D object generation methods across all 2D and 3D metrics by a large margin. These experimental results further demonstrate the superiority of our TAR3D over existing methods.
@@
Triplane Size $\times\!16\!\times\!16 3×3$$\times\!32\!\times\!32$ $3\!\times\!48\!\times\!48$
@@ 4.4 Ablation Studies
[…]
@@ 5 Conclusions & Future Work
In this paper, we propose a novel framework named TAR3D to generate high-quality 3D assets via the “next-token prediction” paradigm derived from multimodal large language models. To achieve this, we first develop a 3D VQ-VAE, in which the 3D shapes are encoded into the triplane latent space and quantized as discrete embeddings from a trainable codebook. Equipped with the well-designed feature deformation and attention mechanism, these quantized features are utilized to reconstruct fine-grained geometries in a neural occupancy field. Then, we adopt the codebook indices of these discrete representations to form the sequence for autoregressive modeling. Driven by the prefilling prompt embeddings, our 3D GPT, along with the 3D spatial information from the proposed TriPE, can generate high-quality 3D objects part by part.
@@
We believe it is promising to improve 3D object generation under the autoregressive setting. There are a few foreseeable directions to explore: 1) Scaling law. We will collect more 3D data with prompts from other datasets, e.g., Objaverse-XL [18], 3D-FUTURE [23], and scale the GPT model. 2) Sequence formulation. We will explore the sequence formulation manner of triplane indices for more efficient generative modeling.
Acknowledgments. This research was partially funded by Shenzhen Science and Technology Program (No. JCYJ20240813114237048).